Q. Two fair coins are tossed. What is the probability of getting at least one Tail?

  • (A) 1/4
  • (B) 1/2
  • (C) 3/4
  • (D) 1
βœ… Correct Answer: (C) 3/4
Explanation: Using complement: P(no Tail) = P(HH) = 1/4. Required probability = 1 - 1/4 = 3/4.
Explanation by: Krishna Sharma
Using complement: P(no Tail) = P(HH) = 1/4. Required probability = 1 - 1/4 = 3/4.

πŸ“ Notes by Other Students/ Teachers


πŸ“Š Question Analytics

πŸ‘οΈ
41
Total Visits
πŸ“½οΈ
4 mo ago
Published
πŸŽ–οΈ
Krishna Sharma
Publisher
πŸ“ˆ
98%
Success Rate